For several years in Massachusetts, the lottery commission would mail residents coupons for free Lotto tickets. To win the jackpot in Massachusetts, you have to correctly guess all six numbers drawn from a pool of 36. What is the expected value of the free Lotto ticket if the jackpot is $8,000,000 and there is no splitting of the prize?
Answer by stanbon(75887)

Massachusetts, you have to correctly guess all six numbers drawn from a pool of 36. What is the expected value of the free Lotto ticket if the jackpot is $8,000,000 and there is no splitting of the prize?
Probability of winning is 1/36C6= 1/1947792
Expected winnings are (1/1947792)(8,000000)=$4.10
Cheers,
Stan H.